# Boosted second moment method in random regular graphs

### v1

@@ 1.1. Labelings and entropy
Later on we will need the Shannon entropy of $\pi$ and $\pi_{\mathrm{vtx}}$ defined as follows:
(2) $H(\pi)=\sum_{x,y\in\mathcal{X}}h\big(p_{x,y}\big)\quad\text{and}\quad H(\pi_{\mathrm{vtx}})=\sum_{x\in\mathcal{X}}h\big(p_{x}\big),\\ \text{where }h(x)=\begin{cases}-x\log x&\text{if }x\in(0,1];\\ 0&\text{if }x=0.\end{cases}$
Many fundamental problems in graph theory can be phrased as questions about whether a graph has a labeling with certain edge statistics. For instance, the independence ratio of $G$ is at least $\alpha$ if and only if $G$ has $\{0,1\}$-labeling with edge statistics
$p_{1,1}=0\quad\text{and}\quad p_{0,1}=p_{1,0}\geq\alpha.$
Recall that $\mathbb{G}_{N,d}$ denotes the uniform random graph among all $d$-regular simple graphs on the vertex set $\{1,\ldots,N\}$. This is closely related to the random multigraph $\mathbb{G}^{\textrm{conf}}_{N,d}$ produced by the so-called configuration model. The exponential rate of the expected number of labelings of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with edge statistics close to a given $\pi$ can be seen to be equal to
@@
A second moment argument (supplemented by an Azuma-type concentration result) gives the following result.
@@
Next we explain how one can boost this result by considering larger neighborhood statistics. We start with defining the notion of tree-indexed Markov chains. Let $T_{d}$ denote the $d$-regular infinite tree. A $T_{d}$-indexed Markov chain is a random labeling $X\colon V(T_{d})\to\mathcal{X}$ with the property that, conditioned on the label $X_{v}$ of a vertex $v\in V(T_{d})$, $X$ is conditionally independent on the $d$ subtrees obtained after deleting $v$. (Similar properties are often referred to as spatial Markov properties, while the authors of [1] call such random labelings vertex-Markov processes.) It is easy to see that the distribution $\mu$ of $X$ is uniquely determined by its marginal on two neighboring vertices, which can be considered as some $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$. For $R\in\mathbb{N}$, let $\mu_{R}$ denote the marginal of $\mu$ corresponding to an $R$-ball of $T_{d}$. It is not hard to show that the following stronger version of the second moment method also holds true; see [1, Thm 1.15].
Claim.
Let $\mu$ be a $T_{d}$-indexed Markov chain with edge marginal $\pi\in\mathcal{P}_{\mathrm{edge}}^{\mathcal{X}}$. If $\pi$ satisfies the second moment condition (3), then, for any $\varepsilon>0$ and any $R\in\mathbb{N}$, $\mathbb{G}_{N,d}$ a.a.s. has an $\mathcal{X}$-labeling with $R$-neighborhood statistics $\varepsilon$-close to $\mu_{R}$.
In essence, the second moment condition guarantees the existence of a labeling that satisfies the spatial Markov property in the sense of local statistics. Therefore, we know what our labeling looks like in larger neighborhoods. Then we can use (randomized) local algorithms to produce new labelings: the new label of a vertex $v$ is computed based on the original labels in a neighborhood of $v$. The point is that we can compute the local statistics of the new labeling using the large-neighborhood statistics of the original one. Next we demonstrate this approach in the context of independent sets.
1.2. Independent sets
We already pointed out that an independent set can be thought of as a $\{0,1\}$-labeling with no adjacent $1$’s. So in the edge statistics of the labeling we have $p_{1,1}=0$, while $p_{0,1}=p_{1,0}=p_{1}$ is the density of the independent set that we usually denote by $\alpha$, and hence $p_{0,0}=1-2\alpha$.
Using the method outlined in Section 1.1, we will find a “Markovian independent set” with some density $\alpha$. Then one can design local algorithms that produce a larger independent set with density $\hat{\alpha}>\alpha$. In this paper we use a very basic algorithm. We call a vertex full-zero if its label is 0 and the labels of all its neighbors are also 0. Consider the induced subgraph over the full-zero vertices. We will see that the connected components are finite with probability $1$. Consequently, one may choose an independent set inside each finite component, and add these to our original independent set. It follows that the original Markovian independent set can be augmented by all one-vertex components and at least half of the remaining full-zero vertices. Since our starting independent set is Markovian, we can easily work out the density $\hat{\alpha}$ of our new independent set by computing the probability that a vertex is full-zero and the probability that it is an isolated full-zero vertex, see Section 4 for details.
One could, of course, consider more sophisticated algorithms to augment the Markovian independent set, resulting in a slightly higher density. However, the improvements would be minor, so we decided to stick with this basic method.
Table 2 shows the largest $\alpha$ that satisfies the second moment condition (3) in the context of independent sets and the corresponding $\hat{\alpha}$ gained by the basic improvement described above (for degrees in the range $3\leq d\leq 11$).
$d$ 3 4 5 6 7 8 9 10 11
$\alpha$ 0.41418 0.36580 0.33268 0.30776 0.28791 0.27151 0.25758 0.24552 0.23491
$\hat{\alpha}$ 0.42718 0.38339 0.35089 0.32525 0.30427 0.28663 0.27152 0.25837 0.24677
Table 2. Density of the Markovian and the augmented independent set
1.3. Further applications
To summarize, our approach is to use a stronger version of the second moment method to find some “typical” Markovian labeling, then perform certain local modifications. Note that the new (possibly non-Markovian) labeling may be a different “type of object” than the original Markovian labeling. In particular, we can start with the same Markovian independent set and create other objects, and hence proving the existence of more complex structures in random regular graphs. (For instance, suppose that we want to find large induced subgraphs with degree bound, say, $2$. It is not hard to design a local algorithm that produces such a subgraph from an independent set.)
In this paper we present one such application, concerning so-called star decompositions. In Section 5 we demonstrate how our approach can be used to prove that, with high probability, the edge set of a random $d$-regular graph can be decomposed into $k$-stars for certain pairs $d,k$.
Organization of the paper
Section 2 gathers the necessary facts about invariant and typical processes on the infinite tree. In Section 3 we examine the second moment condition for independent sets, while in Section 4 we discuss how the Markovian independent set can be augmented. Finally, Section 5 is concerned with star decompositions.
@@ 2.1. Processes
Suppose that $\mathcal{X}$ is a finite set of colors/labels. A mapping $V(G)\to\mathcal{X}$ is a vertex labeling of the graph $G$. A random labeling $X\colon V(T_{d})\to\mathcal{X}$ is a $T_{d}$-indexed Markov chain if, conditioned on the label $X_{v}$ of a vertex $v\in V(T_{d})$, $X$ is conditionally independent on the $d$ subtrees obtained after deleting $v$.
@@
An invariant process $\mu$ is said to be typical [^3] In [1, Def 1.1] the term strongly typical was used. if it can be mimicked on most $d$-regular graphs in the following sense. For every $\varepsilon>0$ and $R\in\mathbb{N}$ there exists $N_{0}=N_{0}(\varepsilon,R)$ such that for every $N>N_{0}$ the following holds for $\mathbb{G}_{N,d}$ with probability at least $1-\varepsilon$: the graph $\mathbb{G}_{N,d}$ has an $\mathcal{X}$-labeling with $R$-neighborhood statistics $\varepsilon$-close [^4] When we talk about the distance of discrete distributions, let us use, say, the total variation distance. to that of $\mu$.
@@ 2.2. Entropy
See (2) for how Shannon entropy $H(\cdot)$ is defined. The motivation for this definition is that $\Sigma(\pi)$ gives the exponential rate of the expected number of $\mathcal{X}$-labelings of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with edge statistics close to $\pi$. In particular, if $\Sigma(\pi)<0$, then $\mathbb{G}^{\textrm{conf}}_{N,d}$ may have such a labeling only with exponentially small probability (in $N$). This fact is what we usually refer to as the first moment method in random regular graphs.
@@
We will use the following stronger form of the second moment method.
@@
We need to consider distributions $\nu$ over $\mathcal{X}^{4}$ that is invariant under exchanging the first and second vertex [^5] Note that $\nu$ does not have to be invariant under exchanging the two colors. and its marginal for any color is $\mu_{\mathrm{edge}}$. We will use the notation
@@
Note that the comma separates the labels of the first and second vertex of the edge. With this notation, the property that $\nu$ is exchangeable can be expressed as
@@
As for the condition on the marginals, we have
@@
Consider the independent coupling $\nu^{\star}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mu_{\mathrm{edge}}\times\mu_{\mathrm{edge}}$ for which $p_{ij,k\ell}=p_{i,k}p_{j,\ell}$ (that is, the red and blue labels are independent). We clearly have
@@
Note that if we choose $\nu$ to be the distribution of two identical copies of $\mu_{\mathrm{edge}}$ (that is, the blue and red labels are always the same), then $\Sigma(\nu)=\Sigma(\mu_{\mathrm{edge}})$. Therefore, if the second moment condition holds, then we necessarily have the “first moment condition” $\Sigma(\mu_{\mathrm{edge}})\geq 0$.
@@ 3. Second moment for independent sets
By the density we mean the probability that a given vertex has label $1$. Note that a Markovian independent set is uniquely determined by its density.
Fix some degree $d\geq 3$ and let $\mu=\mu(d,\alpha)$ denote the distribution of the Markovian independent set on $T_{d}$ with density $\alpha$. Our goal in this section is to prove that $\mu$ is typical for some $\alpha$ (as large as possible).
@@
Let $\beta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{11}$. (This corresponds to the density of the intersection of the blue and red independent sets. Note that this is different from $p_{1,1}=0$, which is the probability that two neighboring vertices have 0 blue labels.) Then we have
@@
Now let $\gamma\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{10,01}=p_{01,10}$. Then
@@
For fixed $\alpha,\beta$, the function $\gamma\mapsto H(\nu)$ has the following derivative:
$2\log\gamma+4\log(\alpha-\beta-\gamma)-2\log(1-4\alpha+2\beta+2\gamma).$
It is easy to see that this is is monotone decreasing in $\gamma\in[0,\alpha-\beta]$ and it vanishes if and only if
$(\alpha-\beta-\gamma)^{2}=\gamma(1-4\alpha+2\beta+2\gamma),$
which leads to
(9) $\gamma=(\alpha-1/2)+\sqrt{(\alpha-1/2)^{2}+(\alpha-\beta)^{2}}.$
Therefore, $H(\nu)$ attains its maximum at (9) for any fixed $\alpha,\beta$.
@@
and show that the maximum is attained at the independent coupling $\nu^{\star}$ (see Remark 2.5). Since $H(\nu_{\mathrm{vtx}})$ does not depend on $\gamma$, we may make the substitution (9) for $\gamma$. We get a function in $\alpha,\beta$. In fact, if we treat $d$ and $\alpha$ as parameters and $\beta$ as the single variable, we get the following function $f_{d,\alpha}(\beta)$ expressing $\max_{\gamma}\Sigma(\nu)$:
@@
The second moment condition is satisfied for a given $\alpha$ if this function attains its maximum at $\beta=\alpha^{2}$ (corresponding to the independent coupling). We summarize our findings in the following theorem.
Theorem 3.2.
Let $d\geq 3$ be an integer and let $\alpha\in(0,1/2)$. Suppose that the function $f_{d,\alpha}\colon[0,\alpha]\to\mathbb{R}$ attains its maximum at $\alpha^{2}$. Then the Markovian independent set of density $\alpha$ is a typical process on $T_{d}$.
It turns out that, when $\alpha$ is in the relevant region, the function $f_{d,\alpha}$ has two local maxima, one at $\alpha^{2}$ and another one above $\alpha^{2}$. (It is actually close to $\alpha$ for large $d$). So one needs to check which local maximum is larger, and find (the largest possible) $\alpha$ for which the global maximum is at $\beta=\alpha^{2}$ (corresponding to the independent coupling). This is quite straightforward to do with the help of a computer.
See Table 2 in the introduction for the largest $\alpha$ satisfying the second moment condition for small degrees.

### v2

@@ 2.2. Labelings and entropy
This section is primarily intended for readers who are unfamiliar with the language of graph limits used in the subsequent sections. Our aim is to explain the relationship between counting in $\mathbb{G}^{\textrm{conf}}_{N,d}$ and entropic quantities in the context of graph labelings. It will shed some light on where the formulas for $\varphi_{d}$ and $f_{d,\alpha}$ come from.
@@
The key quantities here are the Shannon entropy of $\pi$ and $\pi_{\mathrm{vtx}}$ defined as follows:
(11) $H(\pi)=\sum_{x,y\in\mathcal{X}}h\big(p_{x,y}\big)\quad\text{and}\quad H(\pi_{\mathrm{vtx}})=\sum_{x\in\mathcal{X}}h\big(p_{x}\big).$
The exponential rate of the expected number of labelings of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with edge statistics close to a given $\pi$ can be seen to be equal to
@@
For a rigorous proof see [2, Lemma 4.1 and the proof of Theorem 4].
Therefore, the first moment method in this context can be phrased as follows.
Claim.
If $\Sigma(\pi)<0$, then there exists $\varepsilon>0$ such that $\mathbb{G}_{N,d}$ a.a.s. has no $\mathcal{X}$-labeling with edge statistics $\varepsilon$-close to $\pi$.
Furthermore, a second moment argument (supplemented by a standard Azuma-type concentration result) gives the following. (Note that Lemma 3.3 of the next section contains a stronger result that produces labelings satisfying a certain spatial Markov property.)
@@
Many fundamental problems in graph theory can be phrased as questions about whether a graph has a labeling with certain edge statistics. For instance, an independent set of density $\alpha$ corresponds to a $\{0,1\}$-labeling with edge statistics
$p_{1,1}=0\quad\text{and}\quad p_{0,1}=p_{1,0}=\alpha.$
As we will see in Section 4, in this special case we get
$\Sigma(\pi)=\varphi_{d}(\alpha)\quad\text{and}\quad\max\Sigma(\nu)=\max f_{d,\alpha}=\varphi_{d}^{(2)}(\alpha).$
@@ 3.1. Processes
Suppose that $\mathcal{X}$ is a finite set of colors/labels. A mapping $V(G)\to\mathcal{X}$ is a vertex labeling of the graph $G$. A random labeling $X\colon V(T_{d})\to\mathcal{X}$ is a $T_{d}$-indexed Markov chain [^8] Similar properties are often referred to as spatial Markov properties, while the authors of [1] call such random labelings vertex-Markov processes. if, conditioned on the label $X_{v}$ of a vertex $v\in V(T_{d})$, $X$ is conditionally independent on the $d$ subtrees obtained after deleting $v$.
@@
An invariant process $\mu$ is said to be typical [^9] In [1, Def 1.1] the term strongly typical was used. if it can be mimicked on most $d$-regular graphs in the following sense. For every radius $R\in\mathbb{N}$ and every positive $\varepsilon$, there exists $N_{0}=N_{0}(\varepsilon,R)$ such that for each $N>N_{0}$ the following holds for $\mathbb{G}_{N,d}$ with probability at least $1-\varepsilon$: the graph $\mathbb{G}_{N,d}$ has an $\mathcal{X}$-labeling with $R$-neighborhood statistics $\varepsilon$-close [^10] When we talk about the distance of discrete distributions, let us use, say, the total variation distance. to that of $\mu$.
@@ 3.2. Entropy
See (11) for how Shannon entropy $H(\cdot)$ is defined. The motivation for this definition is that $\Sigma(\pi)$ gives the exponential rate of the expected number of $\mathcal{X}$-labelings of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with edge statistics close to $\pi$ (see Section 2.2). In particular, if $\Sigma(\pi)<0$, then $\mathbb{G}^{\textrm{conf}}_{N,d}$ has such a labeling only with exponentially small probability (in $N$). This fact is what we usually refer to as the first moment method in random regular graphs.
@@
We will use the following strong form of the second moment method.
@@
We need to consider distributions $\nu$ over $\mathcal{X}^{4}$ with the following two properties:
  • (i)
    $\nu$ is invariant under exchanging the first and second vertex; [^11] Note that $\nu$ does not have to be invariant under exchanging the two colors.
  • (ii)
    the marginal of $\nu$ corresponding to any of the two colors is $\mu_{\mathrm{edge}}$.
We will use the notation
@@
Note that the comma separates the labels of the first and second vertex of the edge. With this notation, property (i) can be expressed as
@@
As for property (ii) regarding the marginals, it translates to
@@
Consider the independent coupling $\nu^{\star}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mu_{\mathrm{edge}}\times\mu_{\mathrm{edge}}$ for which $p_{ij,k\ell}=p_{i,k}p_{j,\ell}$ (that is, the red and blue labels are independent). This clearly satisfies the above properties and we have
@@
Note that if we choose $\nu$ to be the distribution of two identical copies of $\mu_{\mathrm{edge}}$ (that is, the blue and red labels are always the same), then $\Sigma(\nu)=\Sigma(\mu_{\mathrm{edge}})$. Therefore, if the second moment condition (14) holds, then we necessarily have the “first moment condition” $\Sigma(\mu_{\mathrm{edge}})\geq 0$.
@@ 4. Second moment for independent sets
By its density we mean the probability that a given vertex has label $1$. Note that a Markovian independent set is uniquely determined by its density.
Fix some degree $d\geq 3$ and let $\mu=\mu(d,\alpha)$ denote the distribution of the Markovian independent set on $T_{d}$ with density $\alpha$. Our goal in this section is to prove that $\mu$ is typical provided that $d$ and $\alpha$ satisfy a certain second moment condition.
@@
Let $\beta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{11}$. (This corresponds to the density of the intersection of the blue and red independent sets. Note that this is different from $p_{1,1}=0$, which is the probability that two neighboring vertices both have 0 blue labels.) Then we have
@@
Next let $\gamma\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{10,01}=p_{01,10}$. Then
@@
and show that the maximum is attained at the independent coupling $\nu^{\star}$ (see Remark 3.5). For fixed $d$ and $\alpha$, this expression is a two-variable function in $\beta$ and $\gamma$ that we denote by $\hat{f}_{d,\alpha}$:
(18) $\begin{split}\hat{f}_{d,\alpha}(\beta,\gamma)&\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{d}{2}\bigg(2h(\beta)+2h(\gamma)+4h(\alpha-\beta-\gamma)+h(1-4\alpha+2\beta+2\gamma)\bigg)\\ &-(d-1)\bigg(h(\beta)+2h(\alpha-\beta)+h(1-2\alpha+\beta)\bigg).\end{split}$
Note that second line does not depend on $\gamma$, while the first line is concave in $\gamma$ because $h$ is a concave function. Therefore, for any fixed $\beta$, $\hat{f}_{d,\alpha}(\beta,\gamma)$ is concave in $\gamma$, and takes its maximum, where the partial derivative w.r.t. the second variable vanishes. We have
$\partial_{2}\hat{f}_{d,\alpha}(\beta,\gamma)=d\bigg(-\log\gamma+2\log(\alpha-\beta-\gamma)-\log(1-4\alpha+2\beta+2\gamma)\bigg),$
which vanishes if and only if
$(\alpha-\beta-\gamma)^{2}=\gamma(1-4\alpha+2\beta+2\gamma).$
This is a quadratic equation in $\gamma$ with one positive root that we denote by
(19) $\Gamma(\beta)$ . . $\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-(1/2-\alpha)+\sqrt{(1/2-\alpha)^{2}+(\alpha-\beta)^{2}}$
(20) $=\frac{(\alpha-\beta)^{2}}{(1/2-\alpha)+\sqrt{(1/2-\alpha)^{2}+(\alpha-\beta)^{2}}}.$
In conclusion, $\hat{f}_{d,\alpha}(\beta,\gamma)$ is maximized in the second variable at $\gamma=\Gamma(\beta)$. We define the one-variable function $f_{d,\alpha}$ as follows:
(21) $f_{d,\alpha}(\beta)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\max_{\gamma}\hat{f}_{d,\alpha}(\beta,\gamma)=\hat{f}_{d,\alpha}\big(\beta,\Gamma(\beta)\big).$
There are two important particular substitutions, namely, $\beta=\alpha$, $\gamma=0$ (identical coupling) and $\beta=\gamma=\alpha^{2}$ (independent coupling). Next we show that, in these special cases, the value of $\hat{f}_{d,\alpha}$ can be expressed using the first moment rate-function $\varphi_{d}(\alpha)$ defined in (7).
Claim 4.2.
We have
$\Gamma(\alpha)=0\text{ and }f_{d,\alpha}(\alpha)$ $=\hat{f}_{d,\alpha}(\alpha,0)=\varphi_{d}(\alpha);$
$\Gamma(\alpha^{2})=\alpha^{2}\text{ and }f_{d,\alpha}(\alpha^{2})$ $=\hat{f}_{d,\alpha}(\alpha^{2},\alpha^{2})=2\varphi_{d}(\alpha).$
Proof.
Due to (20) we clearly have $\Gamma(\alpha)=0$, while $\Gamma(\alpha^{2})=\alpha^{2}$ follows from (19) and the identity
(22) $\big(1/2-\alpha\big)^{2}+\big(\alpha-\alpha^{2}\big)^{2}=\big(1/2-\alpha+\alpha^{2}\big)^{2}.$
Then $f_{d,\alpha}(\alpha)=\varphi_{d}(\alpha)$ and $f_{d,\alpha}(\alpha^{2})=2\varphi_{d}(\alpha)$ follow easily from (18) if we make the particular substitutions. ∎
Our aim is to apply Lemma 3.3 in the above setting. Here the independent coupling corresponds to the probabilities
$\beta=p_{1}^{2}=\alpha^{2}\quad\text{and}\quad\gamma=p_{1,0}p_{0,1}=\alpha^{2}.$
Therefore, the condition of the lemma is that $\hat{f}_{d,\alpha}$ is maximized at $\beta=\gamma=\alpha^{2}$. According to Claim 4.2, this is the same as saying that $f_{d,\alpha}$ is maximized at $\beta=\alpha^{2}$, and the following result is obtained.
Theorem 4.3.
Let $d\geq 3$ be an integer and let $\alpha\in(0,1/2)$. If condition (3) holds (that is, $f_{d,\alpha}$ attains its maximum at $\alpha^{2}$), then $\mu(d,\alpha)$, the Markovian independent set of density $\alpha$, is a typical process on $T_{d}$.
Note that, using Definition 1.2, we can simply state this result as follows:
$\mu(d,\alpha^{\mathrm{SM}}_{d})\text{ is typical; in particular, }\alpha^{\star}_{d}\geq\alpha^{\mathrm{SM}}_{d}.$
The question arises: what does the graph of $f_{d,\alpha}$ look like, and how can we check if condition (3) holds for specific $d$ and $\alpha$? In Figure 1 we plotted $f_{d,\alpha}(\beta)$ on $0,\alpha]$ in the case $d=100$ and $\alpha=0.064$. The first local maximum corresponds to $\beta=\alpha^{2}$. Since it is larger than the second local maximum, we can conclude that (3) is satisfied, and hence the Markovian independent set of density $0.064$ is typical on $T_{100}$.
Figure 1. The plot of $f_{d,\alpha}(\beta)$ for $d=100$ and $\alpha=0.064$. The red dot corresponds to $\beta=\alpha^{2}$ (independent coupling). The second moment method works here because the red dot is above the second local maximum (black dot) so it is the global maximum.
Do we always see a similar picture? How does the plot of $f_{d,\alpha}$ change as $\alpha$ varies? We encounter the following behavior [^12] We will not prove all of the claimed behavior; only the pieces necessary for proving Theorem 1.3 are included in the paper. See Section 8 for precise statements and rigorous proofs. for each $d\geq 3$. There are two density thresholds (the second of which can be expressed with an explicit formula):
$0<\alpha_{d}^{\textrm{1st}}<\alpha_{d}^{\textrm{2nd}}=\frac{\sqrt{d-1}-1}{d-2}<\frac{1}{2}.$
There are three phases:
  • $0<\alpha<\alpha_{d}^{\textrm{1st}}$:
    $f_{d,\alpha}$ has a single local maximum at $\alpha^{2}$, which is, therefore, also the global maximum, and hence the second moment condition (3) is always satisfied.
  • $\alpha_{d}^{\textrm{1st}}<\alpha<\alpha_{d}^{\textrm{2nd}}$:
    $f_{d,\alpha}$ has two local maxima: one at $\beta_{1}^{\max}=\alpha^{2}$ and one at some $\beta_{2}^{\max}>\alpha^{2}$. The second moment condition (3) is satisfied if and only if
    (23) $f_{d,\alpha}\big(\beta_{1}^{\max}\big)\geq f_{d,\alpha}\big(\beta_{2}^{\max}\big).$
  • $\alpha_{d}^{\textrm{2nd}}<\alpha<1/2$:
    $f_{d,\alpha}$ has a local minimum at $\beta=\alpha^{2}$, so the second moment condition (3) is never satisfied.
Therefore, in the intermediate phase (between $\alpha_{d}^{\textrm{1st}}$ and $\alpha_{d}^{\textrm{2nd}}$), checking condition (3) amounts to finding $\beta_{2}^{\max}$ and verifying (23). (For specific $d$ and $\alpha$, this can be done numerically with a computer very quickly.)
We finish this section with the following simple observation.
Remark 4.4.
Due to Claim 4.2 we have
@@
Therefore, $\alpha^{\mathrm{SM}}_{d}\leq\alpha^{\mathrm{FM}}_{d}$ (which also follows from the fact that $\alpha^{\star}_{d}$ is sandwiched between the two).